# Spectrum of slip dynamics, scaling & statistical laws emerge from simplified model of fault and damage zone architecture

### v1

@@ Abstract
Seismological and geodetic observations of a fault zone reveal a wide range of slip dynamics, scaling, and statistical laws. However, the underlying physical mechanisms remain unclear. In this study, we show that incorporating an off-fault damage zone-characterized by distributed fractures surrounding a main fault-can reproduce many key features observed in seismic and geodetic data. We model a 2D shear fault zone in which off-fault cracks follow power-law size and density distributions, and are oriented either optimally or parallel to the main fault. All fractures follow the rate-and-state friction law with parameters chosen such that each can host slip instabilities. We do not introduce spatial heterogeneities in the frictional properties of the fault. Using quasi-dynamic boundary integral simulations accelerated by hierarchical matrices, we simulate slip dynamics of this system and analyze the events produced both on and off the main fault. Despite the spatially uniform frictional properties, we observe a natural continuum from slow to fast ruptures, as observed in nature. Our simulations reproduce the Omori law, the inverse Omori law, the Gutenberg-Richter scaling, and the moment-duration scaling. We also observe seismicity localizing toward the main fault when an event is about to nucleate on the main fault. During slow slip events, off-fault seismicity migrates in a pattern resembling a fluid diffusion front, despite the absence of fluids in the model. We also show that tremors, Very Low Frequency Earthquakes (VLFEs), Low Frequency Earthquakes (LFEs), Slow Slip Events (SSEs), and earthquakes (EQs) can all emerge naturally in the ‘digital twin’ framework.
@@
Earthquake faults show complex behavior that ranges from slow creeping movements to fast and destructive ruptures, but their origin is not very clear. We investigated whether including the network of smaller cracks around a main fault—called the “damage zone”—could explain these observations. We created computer models of fault systems that included both the main fault and the surrounding fracture networks. We found that the geometric arrangement of these fractures alone could produce the full spectrum of fault behaviors observed in nature. Our simulations successfully reproduced well-known earthquake patterns, including how aftershocks decay over time, the relationship between earthquake size and frequency, and how seismic activity appears to migrate along fault zones. All major types of fault slip—from barely detectable slow movements to regular earthquakes—emerged naturally from the same model. This research suggests that the complex network of small fractures surrounding major faults plays a crucial role in earthquake behavior. Understanding these damage zones could improve our ability to assess seismic hazards and better predict how fault systems will behave over time.
@@ 1 Introduction
In the brittle upper crust, fault zone displays a large variety of slip dynamics and moment release. Until the discovery of SSEs [79, 58] and tremor [131], faults were thought to either remain locked during the interseismic period or continually creep-the former leading to stick-slip-like release of stored strain energy [30], the latter to continuous strain release [190]. Modern seismological and geodetic observations have significantly advanced our understanding of fault slip behavior, revealing a continuum of deformation modes spanning from regular fast earthquakes to a diverse family of slow earthquakes and/or steady creep. Slow earthquakes includes “seismic” members such as low-frequency earthquakes (LFEs), tectonic tremor, and very low-frequency earthquakes (VLFEs), as well as “geodetic” members—slow slip events (SSEs)—typically classified as short-term (days to weeks) or long-term (months to years) [128]. These phenomena are increasingly recognized as different manifestations of a shared underlying physical process [20, 81]. SSEs, characterized by the gradual release of tectonic stress, have been observed in numerous subduction zones, including Cascadia, Central Ecuador, Guerrero, Hikurangi, Northern Chile, and southwest Japan, as well as along continental plate boundaries such as the San Andreas fault system in California, North Anatolian fault system, Haiyuan fault [113, 58, 158, 57, 90, 198, 169, 164, 182, 122, 50]. These events often co-occur with LFEs and tremor, forming Episodic Tremor and Slip (ETS) sequences, particularly well documented in Cascadia and Nankai [158, 122]. Since their discovery, observations and models have emphasized the potential role of SSEs in the earthquake cycle. Occurring near or within the seismogenic zone, SSEs are thought to influence the initiation or modulation of large earthquakes [158, 178, 132, 48]. Moreover, aseismic slip phenomena are not limited to traditionally aseismic regions. Increasingly, observations show that slow and fast slip events can coexist within the seismogenic zone itself [174, 86, 169, 194]. Aseismic slip has notably been observed prior to several major earthquakes, including the 2011 Tōhoku-Oki [86], 2014 Iquique [169], and 2017 Valparaíso earthquakes [168, 33]. Postseismic aseismic slip in seismic patches has also been documented in several subduction/collision zones [89, 194, 200], further highlighting the complex interaction between seismic and aseismic processes.
@@
Although slip dynamics may appear highly complex across timescales, robust empirical laws continue to emerge from seismological observations. The Gutenberg-Richter law describes the magnitude-frequency distribution of earthquakes, following a power law with a b-value near 1 [73]. The Omori law characterizes the decay rate of aftershocks following a mainshock [197], while an inverse Omori law describes the acceleration of foreshock activity before major events, first noted by [144]. For fast ruptures, seismic moment scales cubically with rupture duration [83, 81]. For slow ruptures, debate remains between a linear scaling [83] and a cubic one [70, 122]. More recently, [97] reported a phenomenon of localization followed by delocalization of deformation before and after major earthquakes, respectively—a pattern observed in southern California for several large events, including the 1992 $M_{w}$ 7.3 Landers, 1999 $M_{w}$ 7.1 Hector Mine, and 2019 $M_{w}$ 7.1 Ridgecrest earthquakes [19]. Furthermore, seismicity has been observed to migrate during slow-slip events, with spatiotemporal features resembling a diffusion front, suggesting that fluid diffusion may govern these dynamics [[, for e.g.]]Danr2024.
@@
The wide spectrum of observed slip behaviors has prompted investigation into various physical mechanisms, with fault friction heterogeneity emerging as the most commonly invoked [30, 173]. Laboratory experiments have consistently shown that frictional properties are central to the transition from stable to unstable slip, particularly near the brittle-ductile transition zone [[, for e.g.]]Leeman2016, Tinti2016, Scuderi2016, Scuderi2017, Leeman2018, Sirorattanakul2024, Pignalberi2024, Yuan2024, Meyer2024, SalazarVsquez2024. Numerical simulations incorporating variable constitutive friction parameters within rate-and-state frameworks have successfully reproduced a spectrum of slip modes, including slow ruptures [206, 110, 16, 126]. Modeling and experimental studies indicate that frictional heterogeneity not only modulates slip dynamics and triggering [14, 60] but also governs complex precursory processes that may culminate in a mainshock [98, 15, 59, 71, 203]. Other mechanisms have also been proposed, including fluid-related effects such as dilatant strengthening [177, 179, 112], spatiotemporal variations in permeability and pore fluid pressure [187, 49, 208, 145, 143], elastic or poroelastic bimaterial effects [77, 1], brittle patches embedded in ductile matrices [8], and transient stress perturbations due to nearby earthquakes [111] or failed nucleation processes [166]. Notably, the earliest model of slow slip events was proposed by [146], who showed that a spring-block system obeying rate-and-state friction and subjected to normal stress perturbations could, under certain conditions, exhibit “aseismic stick-slip” behavior—prior to the observational discovery of SSEs in Cascadia and the Bungo Channel.
@@
However, faults are far more complex than simple, narrow shear zones idealized as frictional surfaces. Geological observations show that, surrounding the fine-grained fault core—where most slip localizes—is a broader damage zone composed of pervasively fractured rock with an intricate three-dimensional geometry across multiple length scales. [183, 43, 26]. This hierarchical geometric complexity ranges from tens of kilometers [[, see, for e.g.,]]Fletcher2014 down to the millimetric scale [188, 124, 64]. Figure 1 illustrates this hierarchical structure of fault systems over a wide range of length scales. Fault slip typically occurs on a localized plane—the fault core—which lies within a damage zone that generally extends several hundreds of meters in width [44, 184, 149, 17, 185, 171, 140, 157, 109]. Various studies have shown that fault geometry—such as roughness [36, 38, 141], bends [99, 155, 162, 142], branches [13, 93, 92, 133, 21, 24, 115], and step-overs [134, 25, 156], play an important role in the dynamics of earthquake ruptures. Recent work has also underlined the key role of active faults networks [84, 42]. Large earthquakes ($M_{w}>7$) often exhibit complex ruptures involving multiple faults [189]. For example, the 1992 Landers earthquake involved the activation of multiple faults during a single seismic event [188, 68], and the 2016 Kaikōura earthquake involved at least 15 faults [100]. More recently, fault system geometry was found to significantly influence the slip distribution of the $M_{w}$ 7.1 Ridgecrest earthquake [125] and the 2023 $M_{w}$ 7.8 Kahramanmaraş, Turkey, earthquake [40]. Furthermore, recent work by [104] suggests that fault slip stability may be controlled by the orientation and complexity of surrounding fault networks, with complex fault systems associated with locked segments that promote stick-slip behavior, whereas simpler geometries favor stable creep. Several studies have also focused on understanding the role of newly created fracture - or dynamic damage - in the propagation of a single dynamic rupture. This damage can significantly affect the high-frequency radiation generated during rupture propagation [195, 137, 116, 138]. Recent findings highlight radiation from such multifracture structures in subduction zones [39]. Moreover, as a dynamic rupture propagates, the energy dissipated by off-fault frictional damage becomes critical to consider [137]. The fracture energy dissipated within the off-fault volume can be substantial—often comparable to the fracture energy on the main fault itself [12, 137, 138]. Measurements of off-fault inelastic deformation using radar and optical imagery, when compared with aftershock sequences, indicate that fault systems exhibiting greater off-fault damage tend to produce relatively higher numbers of aftershocks [123]. Additionally, dynamic damage has been shown to play a crucial role in controlling the transition to supershear rupture [87].
Figure 1: Hierarchical structure of fault systems over a wide range of length scales. (a) Fault map of Southern California [67]. Black lines indicate fault traces. Stars and colored lines indicate the epicenters and rupture traces of historical earthquakes. (b) Fault map and rupture traces (in red) associated with the 1992 Landers earthquake (modified from [188]). (c) Smaller-scale off-fault fracture network [188]. (d) Schematic of fault zone structure, showing a fault core surrounded by damage zones [124]. Figure adapted from [137].
@@
Thus, while fault slip behavior is strongly influenced by the properties of the interface itself, the overall mechanical behavior of faults is equally shaped by the structural complexity of the entire fault zone. Its physical properties evolve over timescales ranging from seconds to millions of years, accommodating displacements from millimeters to tens of kilometers. In particular, dynamic ruptures can induce significant changes in both on-fault and off-fault microstructures and physicochemical properties. These changes, in turn, affects the style and dynamics of fault slip [[, for e.g.]] Andrews2005, collettini2009, niemeijer2010, thomas2014b, Faulkner2006, dor2006a, dor2006b, Mitchell2009, bhat2010a, Okubo2019, Okubo2020. Understanding such processes inevitably calls for a comprehensive approach that considers the fault zone as a whole, rather than focusing solely on the fault core.
@@
Recent studies increasingly focus on understanding the role of complex and irregular geometries throughout the full seismic cycle. [160] incorporated geometrical complexity in numerical models and found that stress interactions between parallel faults can naturally generate slow slip events. Similarly, [38] showed that foreshock sequences can emerge spontaneously when rough fault surfaces are considered. In a related study, [141] modeled a fault network with subsidiary fractures surrounding a primary rough fault and demonstrated the spontaneous emergence of an Omori-like aftershock decay [197]. Incorporating the realistic geometry of northern Cascadia into numerical models enabled the generation of slow slip events that closely match GPS observations in the region [108]. Although an increasing number of studies explore the effects of geometry over multiple seismic cycles, no current model reproduces the full range of observed slip dynamics [205, 85].
@@
In this paper, we develop a fast and reliable model of quasi-dynamic earthquake cycles capable of incorporating complex fault geometries and networks under a rate-and-state friction law. Our approach is accelerated using hierarchical matrices, enabling the simulation of large, complex systems within reasonable computational time (see Appendix A for more details). Our goal is to investigate the role of realistic fault geometry (called fault volume henceforth) in the dynamics of slip events (at the scale of Figure 1d). We propose a simplified model of a 2D fault volume consisting of a main self-similar rough fault surrounded by a hierarchy of off-fault slip planes. All fractures are frictionally homogeneous (rate-weakening) and capable of dynamic slip. Remarkably, this purely geometric complexity is sufficient to reproduce the full spectrum of observed slip behaviors. The model can generate a continuum of slip events and recover all major empirical scaling laws associated with the seismic cycle.
@@
In the following section, we introduce the fault volume model, stress loading, and initial conditions used in our study. We then outline the methodology for event detection, distinguishing between fast and slow ruptures, and describe our approach for constructing the synthetic earthquake catalog. In the results section, we begin with a case study illustrating the spatiotemporal evolution of seismic cycles and the emergence of complex slip dynamics. We then demonstrate the model’s ability to reproduce key empirical laws, including the Gutenberg-Richter relation, moment-rupture area scaling, moment-duration scaling, and both Omori and inverse Omori laws. Additionally, we highlight the model’s capacity to replicate the localization-delocalization transition of deformation. We further investigate the influence of fracture orientation and frictional behavior on seismic activity. Finally, we present a comprehensive synthetic catalog compiled from all simulations, offering insights into the statistical behavior of our fault volume model.

### v2

# Fault volume digital twin to reproduce the full slip spectrum, scaling and statistical laws

@@ Abstract
Seismological and geodetic observations of fault zones reveal diverse slip dynamics, scaling, and statistical laws. Existing mechanisms explain some but not all of these behaviors. We show that incorporating an off-fault damage zone—characterized by distributed fractures surrounding a main fault—can reproduce many key features observed in seismic and geodetic data. We model a 2D shear fault zone in which off-fault cracks follow power-law size and density distributions, and are oriented either optimally or parallel to the main fault. All fractures follow rate-and-state friction with parameters enabling slip instabilities. We do not introduce spatial heterogeneities in frictional properties. Using quasi-dynamic boundary integral simulations accelerated by hierarchical matrices, we simulate slip dynamics and analyze events produced both on and off the main fault. Despite spatially uniform frictional properties, we observe a natural continuum from slow to fast ruptures, as seen in nature. Our simulations reproduce the Omori law, inverse Omori law, Gutenberg-Richter scaling, and moment-duration scaling. We observe seismicity localizing toward the main fault before nucleation of main-fault events. During slow slip events, off-fault seismicity migrates in patterns resembling fluid diffusion fronts, despite the absence of fluids. We show that tremors, Very Low Frequency Earthquakes, Low Frequency Earthquakes, Slow Slip Events, and earthquakes can all emerge naturally within this fault volume framework, making it an ideal digital twin for testing hypotheses, performing ground-truth inversions, and probing mechanical properties inaccessible with natural observations.
@@
Earthquake faults exhibit complex behavior ranging from slow creeping movements to fast destructive ruptures, but the physical mechanisms underlying this remain unclear. We investigate whether the geometric arrangement of smaller fractures around a main fault—called the “damage zone”—can explain these diverse slip behaviors. We develop simulations that incorporate both a rough main fault and surrounding networks of fractures, all governed by the same frictional properties. We find that geometric complexity alone reproduces the full spectrum of fault slip observed in nature—from barely detectable slow slip events to regular earthquakes—without requiring spatially varying friction. Our simulations naturally generate realistic aftershock sequences following Omori’s law, earthquake size-frequency distributions matching the Gutenberg-Richter relation, spatiotemporal clustering of seismicity, and complex source time functions resembling tectonic tremor. This fault volume model serves as a digital twin that can generate physically consistent synthetic earthquake catalogs for training machine learning algorithms, testing how faults respond to external perturbations like tidal stresses or fluid injection, and developing improved methods for interpreting seismic and geodetic observations. Our results demonstrate that the network of small fractures surrounding major faults provides a fundamental organizing principle for fault dynamics, offering a geometric foundation upon which frictional and hydraulic effects can be evaluated.
Key Points
  • A fault volume model reproduces the full spectrum of slip behaviors—from slow slip events to fast earthquakes—using geometric complexity alone.
  • Damage zone fractures surrounding a rough fault generate realistic aftershock sequences, foreshock activity, and spatiotemporal clustering without imposed frictional heterogeneity.
  • The model naturally recovers empirical scaling laws including Gutenberg-Richter statistics, Omori decay, moment-duration scaling, and source-time-function complexity.
  • Geometric complexity provides a foundational baseline for fault slip dynamics, enabling a digital twin framework for testing physical mechanisms and developing inverse methods.
@@ 1 Introduction
In the brittle upper crust, fault zones display a large variety of slip dynamics and moment release. Until the discovery of slow slip events, SSEs, [70, 50] and tremor [123], faults were thought to either remain locked during the interseismic period or continually creep—the former leading to stick-slip-like release of stored strain energy [26], the latter to continuous strain release [177]. Modern seismological and geodetic observations have significantly advanced our understanding of fault slip behavior, revealing a continuum of deformation modes spanning from regular fast earthquakes to a diverse family of slow earthquakes and/or steady creep. Slow earthquakes include “seismic” members such as low-frequency earthquakes (LFEs), tectonic tremor, and very low-frequency earthquakes (VLFEs), as well as “geodetic” members—slow slip events (SSEs)—typically classified as short-term (days to weeks) or long-term (months to years) [120]. These phenomena are increasingly recognized as different manifestations of a shared underlying physical process [18, 75]. SSEs, characterized by the gradual release of tectonic stress, have been observed in numerous subduction zones, including Cascadia, Central Ecuador, Guerrero, Hikurangi, Northern Chile, and southwest Japan, as well as along continental plate boundaries such as the San Andreas fault system in California, North Anatolian fault system, Haiyuan fault [105, 50, 148, 49, 83, 185, 159, 154, 170, 115, 45]. These events often co-occur with LFEs and tremor, forming Episodic Tremor and Slip (ETS) sequences, particularly well documented in Cascadia and Nankai [148, 115]. Since their discovery, observations and models have emphasized the potential role of SSEs in the earthquake cycle. Occurring near or within the seismogenic zone, SSEs are thought to influence the initiation or modulation of large earthquakes [148, 166, 122, 43]. Moreover, aseismic slip phenomena are not limited to traditionally aseismic regions. Increasingly, observations show that slow and fast slip events can coexist within the seismogenic zone itself [163, 79, 159, 183]. Aseismic slip has notably been observed prior to several major earthquakes, including the 2011 Tōhoku-Oki [79], 2014 Iquique [159], and 2017 Valparaíso earthquakes [158, 29]. Postseismic aseismic slip in seismic patches has also been documented in several subduction/collision zones [82, 183, 188], further highlighting the complex interaction between seismic and aseismic processes.
@@
Although slip dynamics may appear highly complex across timescales, robust empirical laws continue to emerge from seismological observations of fast events. The Gutenberg-Richter law describes the magnitude-frequency distribution of earthquakes, following a power law with a b-value near 1 [64]. The Omori law characterizes the decay rate of aftershocks following a mainshock [184], while an inverse Omori law describes the acceleration of foreshock activity before major events, first noted by [135]. For fast ruptures, seismic moment scales cubically with rupture duration [76, 75]. For slow ruptures, debate remains between a linear scaling [76] and a cubic one [62, 115]. More recently, [90] reported a phenomenon of localization followed by delocalization of deformation before and after major earthquakes, respectively—a pattern observed in southern California for several large events, including the 1992 $M_{w}$ 7.3 Landers, 1999 $M_{w}$ 7.1 Hector Mine, and 2019 $M_{w}$ 7.1 Ridgecrest earthquakes [17]. Furthermore, seismicity has been observed to migrate during slow-slip events, with spatiotemporal features resembling a diffusion front, suggesting that fluid diffusion may govern these dynamics [[, e.g.]]Danr2024.
@@
The wide spectrum of observed slip behaviors has prompted investigation into various physical mechanisms, with fault friction heterogeneity emerging as the most commonly invoked [26, 162]. Laboratory experiments have consistently shown that frictional properties are central to the transition from stable to unstable slip, particularly near the brittle-ductile transition zone [[, e.g.]]Leeman2016, Tinti2016, Scuderi2016, Scuderi2017, Leeman2018, Sirorattanakul2024, Pignalberi2024, Yuan2024, Meyer2024, SalazarVsquez2024. Numerical simulations incorporating variable constitutive friction parameters within rate-and-state frameworks have successfully reproduced a spectrum of slip modes, including slow ruptures [195, 102, 14, 119]. Modeling and experimental studies indicate that frictional heterogeneity not only modulates slip dynamics and triggering [12, 51] but also governs complex precursory processes that may culminate in a mainshock [91, 13, 52, 63, 190]. Other mechanisms have also been proposed, including fluid-related effects such as dilatant strengthening [165, 167, 104], spatiotemporal variations in permeability and pore fluid pressure [174, 44, 197, 136, 134], elastic or poroelastic bimaterial effects [67, 1], brittle patches embedded in ductile matrices [7], and transient stress perturbations due to nearby earthquakes [103] or failed nucleation processes [156]. Notably, the earliest model of slow slip events was proposed by [137], who showed that a spring-block system obeying rate-and-state friction and subjected to normal traction perturbations could, under certain conditions, exhibit “aseismic stick-slip” behavior—prior to the observational discovery of SSEs in Cascadia and the Bungo Channel. All of the above mechanisms primarily focus on modeling the fault as a single frictional surface.
@@
However, faults are far more complex than simple, narrow shear zones idealized as frictional surfaces. Geological observations show that, surrounding the fine-grained fault core—where most slip localizes—is a broader damage zone composed of pervasively fractured rock with an intricate three-dimensional geometry across multiple length scales [171, 39, 23]. This hierarchical geometric complexity ranges from tens of kilometers [[, see, for e.g.,]]Fletcher2014 down to the millimetric scale [175, 117, 57]. Figure 1 of [128] illustrates this hierarchical structure of fault systems over a wide range of length scales. Fault slip typically occurs on a localized plane—the fault core—which lies within a damage zone that generally extends several hundreds of meters in width [40, 172, 139, 15, 173, 160, 131, 147, 101]. Various studies have shown that fault geometry—such as roughness [32, 34, 133], bends [92, 145, 151, 132], branches [11, 86, 85, 124, 19, 21, 108, 180], and step-overs [125, 22, 146], play an important role in the dynamics of earthquake ruptures. Recent work has also underlined the key role of active faults networks [77, 38]. Large earthquakes ($M_{w}>7$) often exhibit complex ruptures involving multiple faults [176]. For example, the 1992 Landers earthquake involved the activation of multiple faults during a single seismic event [175, 60], and the 2016 Kaikōura earthquake involved at least 15 faults [93]. More recently, fault system geometry was found to significantly influence the slip distribution of the $M_{w}$ 7.1 Ridgecrest earthquake [118] and the 2023 $M_{w}$ 7.8 Kahramanmaraş, Turkey, earthquake [36, 193]. Furthermore, recent work by [97] suggests that fault slip stability may be controlled by the orientation and complexity of surrounding fault networks, with complex fault systems associated with locked segments that promote stick-slip behavior, whereas simpler geometries favor stable creep.
Several studies have also focused on understanding the role of newly created fracture - or dynamic damage - in the propagation of a single dynamic rupture [[]among others]xu2015,bhat2012. This damage can significantly affect the high-frequency radiation generated during rupture propagation [182, 128, 109, 129]. Recent findings highlight radiation from such multifracture structures in subduction zones [35], and the anisotropy imposed by the preferred orientation of off-fault fractures [73]. Moreover, as a dynamic rupture propagates, the energy dissipated by off-fault fracturing processes (creation of new fractures or reactivation of preexisting ones) becomes critical to consider [128]. The fracture energy dissipated within the off-fault volume can be substantial—often comparable to the fracture energy on the main fault itself [10, 128, 129]. Measurements of off-fault inelastic deformation using radar and optical imagery, when compared with aftershock sequences, indicate that fault systems exhibiting greater off-fault damage tend to produce relatively higher numbers of aftershocks [116]. Additionally, dynamic damage has been shown to play a crucial role in controlling the transition to supershear rupture [80].
@@
Thus, while fault slip behavior is strongly influenced by the properties of the interface itself, the overall mechanical behavior of faults is equally shaped by the structural complexity of the entire fault zone. Its physical properties evolve over timescales ranging from seconds to millions of years, accommodating displacements from millimeters to tens of kilometers. In particular, dynamic ruptures can induce significant changes in both on-fault and off-fault mechanical properties. These changes, in turn, influence rupture nucleation, propagation, timing, seismic wave radiation, and postseismic deformation. This broader structure alters the rheology of both the fault core and the surrounding rock, introduces complex interactions, and thereby affects the style and dynamics of fault slip [[, e.g.]] Andrews2005, collettini2009, niemeijer2010, thomas2014b, Faulkner2006, dor2006a, dor2006b, Mitchell2009, bhat2010a, biegel2010, bhat2012, Okubo2019, Okubo2020. Understanding such processes inevitably calls for a comprehensive approach that considers the fault zone as a whole, rather than focusing solely on the fault core.
@@
Keeping this in mind, recent studies increasingly focus on understanding the role of complex and irregular geometries throughout the full seismic cycle. [149] incorporated geometrical complexity in numerical models and found that stress interactions between parallel faults can naturally generate slow slip events. Similarly, [34] showed that foreshock sequences can emerge spontaneously when rough fault surfaces are considered. In a related study, [133] modeled a fault network with subsidiary fractures surrounding a primary rough fault and demonstrated the spontaneous emergence of an Omori-like aftershock decay [184]. Incorporating the realistic geometry of northern Cascadia into numerical models enabled the generation of slow slip events that closely match GPS observations in the region [100]. Although an increasing number of studies [194, 78, 196] explore the effects of geometry over multiple seismic cycles, no current model reproduces the full range of observed slip dynamics.
@@
We aim to investigate the role of this realistic fault geometry and its associated off-fault damage—hereafter referred to as the fault volume—in governing slip event dynamics. We propose a simplified model of a 2D fault volume consisting of a main self-similar rough fault surrounded by a hierarchy of off-fault slip planes. All fractures are frictionally homogeneous (rate-weakening) and capable of dynamic slip. To enable simulation of this geometrically complex system, we develop a fast quasi-dynamic earthquake cycle model accelerated using hierarchical matrices. Hierarchical matrices compress the stress interaction matrix by approximating distant interactions with low-rank blocks, enabling simulations of large, complex systems within reasonable computational time (see Supporting Information for more details). Remarkably, this purely geometric complexity is sufficient to reproduce the full spectrum of observed slip behaviors. The model can generate a continuum of slip events and recover all major empirical scaling laws associated with the seismic cycle. This establishes the model framework as a “digital twin” of a fault zone where hypotheses about fault slip mechanisms can be tested systematically, inversion techniques can be validated against known ground-truth, and fundamental mechanical properties—such as stress evolution, energy partitioning, and damage zone dynamics—can be probed directly in ways that are impossible with natural observations.
@@
In the following section, we introduce the fault volume model, stress loading, and initial conditions used in our study. We then outline the methodology for event detection, distinguishing between fast and slow ruptures, and describe our approach for constructing the synthetic earthquake catalog. In the results section, we begin with a case study illustrating the spatiotemporal evolution of seismic cycles and the emergence of complex slip dynamics. We analyze the moment rate functions of events across scales, revealing distinct characteristics between slow and fast ruptures. We then demonstrate the model’s ability to reproduce key empirical laws, including the Gutenberg-Richter relation, moment-rupture area scaling, moment-duration scaling, and both Omori and inverse Omori laws. Additionally, we highlight the model’s capacity to replicate the localization-delocalization transition of deformation and the apparent diffusion-like migration of seismicity during slow slip events. We conclude by discussing the model’s implications and limitations, identifying testable hypotheses, and evaluating its potential as a digital twin framework for fault systems.